Find the linear equation of a line that contains a point of 2, -3 with the slope of 5
Answer by Alan3354(69443)

Find the linear equation of a line that contains a point of 2, -3 with the slope of 5
--------------------------------------
Do it like this one:
==========================
1. through:(3,2), slope=2
y-y1 = m*(x-x1) where (x1,y1) is the point
---
y-2 = 2(x-3)
y = 2x - 4